# On the spectrum of Sturmian Hamiltonians of bounded type in a small coupling regime

### v1

@@ 3.2 Translation to a Toral Map Problem
From Remark 5, there is a $r_{0}=r_{0}\left(N^{*},C^{**},K\right)$ such that
@@
Referring to Lemma 5, let us choose $\rho>0$ so small that $r(\rho)<\frac{r_{0}}{2}$ and set $\lambda_{1}:=\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}$. We see that for $\lambda\in[0,\lambda_{1}]$, the set
$\left\{\boldsymbol{x}\in L_{\lambda}\cap\mathbb{S}_{\lambda,O_{\rho}}:T_{a_{n}}\circ\cdots\circ T_{a_{1}}(\boldsymbol{x})\not\in O_{\rho}\ \text{for all}\ n>0\right\}$ (3.9)
is a one-dimensional compact set that is contained in $L_{\lambda}\cap\Lambda_{\lambda}^{\text{bnd}}$.
It is a straightforward calculation to check that the line $\hat{L}_{0}:=\left[0,\frac{1}{2}\right]\times\{0\}\subset\mathbb{T}^{2}$ is mapped into $L_{0}\cap\mathbb{S}$ under the map $F$. Also, the projection map $\pi_{\lambda}$ sends $L_{\lambda}\cap\mathbb{S}_{\lambda,O_{\rho}}$ onto a curve in $\mathbb{S}$ that is $C^{2}$ close to a segment of $L_{0}$, provided $\lambda$ is small. That is, for small $\lambda$, there is a curve $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus O_{r(\rho)}$, that is $C^{2}$ close to a segment of $\hat{L}_{0}$, that is mapped into $L_{\lambda}$ under the map $F_{\lambda}$, and this curve tends to $\hat{L}_{0}\setminus O_{r(\rho)}$ as $\lambda\rightarrow 0$.
As the map $F_{\lambda}$ preserves Hausdorff dimension, it suffices to estimate the Hausdorff dimension of the set
$\mathcal{C}\left(\lambda,O_{r(\rho)}\right):=\left\{P\in\hat{L}_{\lambda}:\tilde{T}_{\lambda,a_{n}}\circ\cdots\circ\tilde{T}_{\lambda,a_{1}}(P)\not\in O_{r(\rho)}\ \text{for all}\ n>0\right\}.$ (3.10)
That is, it suffices to show that there is a $\lambda_{0}\in(0,\lambda_{1})$ such that
$\text{dim}_{H}\left(\mathcal{C}\left(\lambda,O_{r(\rho)}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ (3.11)
@@
3.3 Extended Torus Maps
@@
where $\tilde{T}_{0,a}^{1}(x,y)=ax+y$, $\tilde{T}_{0,a}^{2}(x,y)=x$, and $\tilde{T}_{\lambda,a}^{1}(x,y),\ \tilde{T}_{\lambda,a}^{2}(x,y):\mathbb{T}^{2}\setminus O_{r(\rho)}\rightarrow S^{1}$, we set
@@
Remark 7.
@@
Since $F\left(\frac{1}{4},\frac{1}{4}\right)=(-1,0,0)$, and $\left(\frac{1}{4},\frac{1}{4}\right)\in\mathbb{T}^{2}\setminus O_{r_{0}}$ (see 3.8), this gives
@@
for any $n\in\mathbb{N}$, and this non-stationary orbit is a finite set. It is similar to check that
$\hat{T}_{\lambda,a_{n+m}}\circ\cdots\circ\hat{T}_{\lambda,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)=\hat{T}_{0,a_{n+m}}\circ\cdots\circ\hat{T}_{0,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)$
for any $m\geq 0$ and $n\in\mathbb{N}$.
@@
The first and second part can be taken as a consequence of Lemma 8. The third part comes from the fact that $A_{a}$ has a positive unstable eigenvalue for each $a$, so if $\hat{T}_{\lambda,a}$ is sufficiently $C^{1}$-close to $\hat{T}_{0,a}$, then the differential of $\hat{T}_{\lambda,a}$ will preserve orientation of vectors from the cone $K$. The last part comes from the fact that the curve $\hat{L}_{\lambda}$ tends to $\hat{L}_{0}\setminus O_{r(\rho)}$ in the $C^{2}$-sense.
From these cone conditions, we have
@@ 3.3.1 Bounded Distortion Estimate
For each $\lambda\in[0,\lambda_{2}]$, if $\gamma$ is a sub-curve of $\hat{L}_{\lambda}$ that has endpoints $P$ and $Q$ and satisfies $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$ for some $i=1,\dots,4$, then for any point $Q^{\prime}\in\gamma\setminus\{P,Q\}$, we must have
@@
$\leq C_{4}\frac{\text{dist}_{\gamma}(P,Q^{\prime})}{\text{dist}_{\gamma}(Q,Q^{\prime})},$
for any $m\geq 0$.
@@
Let $\lambda\in[0,\lambda_{2}]$ and $m\geq 0$. Since $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$, it must have length less than $C_{1}\frac{\sqrt{2}}{10}$. Thus, from Lemma 9 and recalling (3.3), it must be the case that for each $k$ with $0\leq k\leq m$, we have
@@
In addition, since we have that $\hat{L}_{\lambda}$ is a smooth curve tangent to $\mathcal{K}^{u}$ and $\left\|\hat{T}_{\lambda,a_{n}}\right\|\leq C_{2}+1$ for all $n\in\mathbb{N}$, it follows from Remark 6 that
$\sum_{k=0}^{m}\text{Ang}\left(f_{\lambda,k,0}\right)<\Theta.$
@@
3.4 Non-Stationary Stable Manifolds
In this section, we establish a non-stationary version of the Stable Manifold Theorem for our collection of maps.
@@ 3.4 Non-Stationary Stable Manifolds
$W^{s}_{\lambda}\left(Q,m\right):=\left\{P\in\mathbb{T}^{2}:\lim\limits_{n\rightarrow\infty}\text{dist}\left(f_{\lambda,n,m}(P),f_{\lambda,n,m}(Q)\right)=0\right\},$
@@
    $\text{dist}\left(W^{s}_{0,\text{loc}}(Q,k),W^{s}_{\lambda,\text{loc}}(Q,k)\right)<\epsilon$
@@
We prove the result first for $Q=0$. Recalling our notations from Section 2.3.1, we work locally in the neighborhood $\mathbb{B}_{1}=\pi\left(\text{Box}_{\frac{1}{10}}(0)\right)$ and in the coordinate system $\tilde{V}^{\perp}\oplus\tilde{V}$ where $\tilde{V}=\pi\left(V\cap\text{Box}_{\frac{1}{10}}(0)\right)$
@@
From Remark 2 and the cone conditions, we may choose a $\lambda_{3}>0$ such that for each $\lambda\in[0,\lambda_{3}]$ and each $a\in\mathcal{A}$, if $\text{graph}(\varphi)\in X$, then its image under $\hat{T}_{\lambda,a}^{-1}$ must overflow $\mathbb{B}_{1}$ and be tangent to $\mathcal{K}^{s}$.
@@
Through the common cone conditions (Property (C)), for each $\lambda\in[0,\lambda_{3}]$, we find that $\left(\Gamma_{\lambda,a_{n}}\right)$ is a sequence of contractions with uniform contraction rate over $n$. Thus, by Proposition 1, for each $k\in\mathbb{N}$, there is a $\varphi_{\lambda,k}^{*}:V^{\perp}\rightarrow V$ such that $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)\in X$ and
@@
Hence, by Proposition 1, we can make the corresponding fixed points $\text{graph}\left(\varphi_{0,k}^{*}\right)$ and $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)$ arbitrary close in the metric $d_{X}$ for sufficiently small $\lambda$, and this closeness can be made uniform in $k$.
@@
$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq n_{0}\\ m\in\mathbb{N}\end{subarray}}\text{dist}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)<\epsilon,$
@@ 3.5 Rectangle Construction
Recall that by Remark 7, for $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$, we have that
@@
for each $n,m\geq 0$ and $\lambda\in[0,\lambda_{2}]$, and it is also a calculation to see that $\text{Orb}_{0}(Q_{0})$ is a finite set and bounded away from each $Q_{i}$ by at least $\frac{1}{4}$.
@@
This dynamical behavior will allow us to construct a Cantor set contained in $\mathcal{C}\left(\lambda,O_{r(\rho)}\right)$ that has large thickness.
@@ 3.5.1 Base Set of Rectangles
Recalling the constant $C^{*}$ from (3.2) and $C^{**}$ from Lemma 7, we have
@@
    $\partial^{s}R_{m}(\lambda,Q_{i})\subset W^{s}_{\lambda}(Q_{0},m)$ and in particular, +$\partial^{s}_{+}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{+}(Q_{i})$ and $\partial^{s}_{-}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{-}(Q_{i})$
@@
    For any $P\in\partial^{s}R_{m}(\lambda,Q_{i})$, we have $|P-Q_{i}|<\frac{C^{*}}{2}$
@@
We know that $W^{s}_{0}(Q_{0},m)$ is a line through $Q_{0}$ of slope $-\frac{1}{[a_{m},a_{m+1},\dots]}$ and we know from Lemma 9 that
@@
Notice that $f_{0,\tilde{n},m}(P)\in W^{s}_{0,\text{loc}}(Q_{0},m+\tilde{n}+1)$, and so if $n\geq\tilde{n}$, then $f_{\lambda,n,m}(P)\in B_{\frac{1}{10}}(Q_{0})$ so that its distance from $\bigcup_{i=1}^{4}\{Q_{i}\}$ is greater than $\frac{3}{20}\frac{1}{4}-\frac{1}{10}$.
@@
From Corollary 1, the quantity
$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq\tilde{n}\\ m\in\mathbb{N}\end{subarray}}\text{dist}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)$ (3.17)

### v2

@@ 4.3 Choice of Neighborhood
From Proposition 2, there is a $r_{0}=r_{0}\left(N^{*},C^{**},K\right)$ such that
@@
Referring to Lemma 5, let us choose $\rho>0$ so small that $r(\rho)<\frac{r_{0}}{2}$ and set $\lambda_{1}:=\min_{a\in\mathcal{A}}\left\{\tilde{\lambda}(\rho,a)\right\}$.
For $\lambda\in(0,\lambda_{1})$, it is a straightforward calculation to check that the line $\hat{L}_{0}:=\left[0,\frac{1}{2}\right]\times\{0\}\subset\mathbb{T}^{2}$ is mapped into $L_{0}\cap\mathbb{S}$ under the map $F$. Also, the projection map $\pi_{\lambda}$ sends $L_{\lambda}\cap\mathbb{S}_{\lambda,O_{\rho}}$ onto a curve in $\mathbb{S}$ that is $C^{2}$ close to a segment of $L_{0}$, provided $\lambda$ is small. That is, for small $\lambda$, the curve $\hat{L}_{\lambda}\subset\mathbb{T}^{2}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}$ is $C^{2}$ close to a segment of $\hat{L}_{0}$, and in particular, $\hat{L}_{0}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}$ as $\lambda\rightarrow 0$ in the $C^{2}$-sense.
So, to complete the proof of Theorem 4, we will show that there is a $\lambda_{0}\in(0,\lambda_{1})$ such that
$\text{dim}_{H}\left(\mathcal{C}\left(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right)\right)\geq\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$ (4.9)
@@
4.4 Extended Toral Maps
@@
where $\tilde{T}_{0,a}^{1}(x,y)=ax+y$, $\tilde{T}_{0,a}^{2}(x,y)=x$, and $\tilde{T}_{\lambda,a}^{1}(x,y),\ \tilde{T}_{\lambda,a}^{2}(x,y):\mathbb{T}^{2}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\rightarrow S^{1}$, we set
@@
Lemma 12.
For each $\lambda\in\left[0,\lambda_{1}\right]$, we have
$\hat{T}_{\lambda,a_{n+m}}\circ\cdots\circ\hat{T}_{\lambda,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)=\hat{T}_{0,a_{n+m}}\circ\cdots\circ\hat{T}_{0,a_{m}}\left(\frac{1}{4},\frac{1}{4}\right)$ (4.10)
for any $m\geq 0$ and $n\in\mathbb{N}$.
Proof.
@@
Since $F\left(\frac{1}{4},\frac{1}{4}\right)=(-1,0,0)$, and $\left(\frac{1}{4},\frac{1}{4}\right)\in\mathbb{T}^{2}\setminus O_{r_{0}}$ (see 4.8), this gives
@@
for any $n\in\mathbb{N}$, and this non-stationary orbit is a finite set. This same argument gives (4.10) for any $m\geq 0$. ∎
4.4.1 Further Choice of Constants.
@@
The first and second part can be taken as a consequence of Lemma 11. The third part comes from the fact that $A_{a}$ has a positive unstable eigenvalue for each $a$, so if $\hat{T}_{\lambda,a}$ is sufficiently $C^{1}$-close to $\hat{T}_{0,a}$, then the differential of $\hat{T}_{\lambda,a}$ will preserve orientation of vectors from the cone $K$. The last part comes from the fact that the curve $\hat{L}_{\lambda}$ tends to $\hat{L}_{0}\setminus{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}$ in the $C^{2}$-sense.
4.5 Properties of the Extended Toral Maps
This subsection is dedicated to deriving useful dynamical properties of our extended maps $\left\{\hat{T}_{\lambda,a}\right\}$.
4.5.1 Lengths and Orientation of Curves
From Property (C), we have
@@ 4.5.2 Bounded Distortion Estimate
For each $\lambda\in[0,\lambda_{2}]$, if $\gamma$ is a sub-curve of $\hat{L}_{\lambda}$ that has endpoints $P$ and $Q$ and satisfies $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$ for some $i=1,\dots,4$ and $m\geq 0$, then for any point $Q^{\prime}\in\gamma\setminus\{P,Q\}$, we must have
@@
$\leq C_{4}\frac{\text{dist}_{\gamma}(P,Q^{\prime})}{\text{dist}_{\gamma}(Q,Q^{\prime})}.$
@@
Let $\lambda\in[0,\lambda_{2}]$. If $f_{\lambda,m,0}(\gamma)\in\mathbb{B}_{i}$, it must have length less than $C_{1}\frac{\sqrt{2}}{10}$. Thus, from Lemma 13 and recalling (4.3), it must be the case that for each $k$ with $0\leq k\leq m$, we have
@@
In addition, since we have that $\hat{L}_{\lambda}$ is a smooth curve tangent to $\mathcal{K}^{u}$ and $\left\|\hat{T}_{\lambda,a_{n}}\right\|\leq C_{2}+1$ for all $n\in\mathbb{N}$, it follows from Lemma 8 that
$\sum_{k=0}^{m}\text{Ang}\left(f_{\lambda,k,0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\gamma\right)}\right)<\Theta.$
@@
4.5.3 Non-Stationary Stable Manifolds
In this subsection, we establish a non-stationary version of the Stable Manifold Theorem for our collection of maps.
@@ 4.5.3 Non-Stationary Stable Manifolds
$W^{s}_{\lambda}\left(Q,m\right):=\left\{P\in\mathbb{T}^{2}:\lim\limits_{n\rightarrow\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left|f_{\lambda,n,m}(P)-f_{\lambda,n,m}(Q)\right|}=0\right\},$
@@
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{dist}_{H}}\left(W^{s}_{0,\text{loc}}(Q,k),W^{s}_{\lambda,\text{loc}}(Q,k)\right)<\epsilon$
@@
We prove the result first for $Q=0$. Recalling our notations from Section 3.1, we work locally in the neighborhood $\mathbb{B}_{1}=\pi\left(\text{Box}_{\frac{1}{10}}(0)\right)$ and in the coordinate system $\tilde{V}^{\perp}\oplus\tilde{V}$ where $\tilde{V}=\pi\left(V\cap\text{Box}_{\frac{1}{10}}(0)\right)$
@@
From Lemma 6 and the cone conditions, we may choose a $\lambda_{3}>0$ such that for each $\lambda\in[0,\lambda_{3}]$ and each $a\in\mathcal{A}$, if $\text{graph}(\varphi)\in X$, then its image under $\hat{T}_{\lambda,a}^{-1}$ must overflow $\mathbb{B}_{1}$ and be tangent to $\mathcal{K}^{s}$.
@@
Through the common cone conditions (Property (C)), for each $\lambda\in[0,\lambda_{3}]$, we find that $\left(\Gamma_{\lambda,a_{n}}\right)$ is a sequence of contractions with uniform contraction rate over $n$. Thus, by Proposition A.1, for each $k\in\mathbb{N}$, there is a $\varphi_{\lambda,k}^{*}:V^{\perp}\rightarrow V$ such that $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)\in X$ and
@@
Hence, by Proposition A.1, we can make the corresponding fixed points $\text{graph}\left(\varphi_{0,k}^{*}\right)$ and $\text{graph}\left(\varphi_{\lambda,k}^{*}\right)$ arbitrary close in the metric $d_{X}$ for sufficiently small $\lambda$, and this closeness can be made uniform in $k$.
@@
$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq n_{0}\\ m\in\mathbb{N}\end{subarray}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{dist}_{H}}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)<\epsilon,$
@@ 4.6 Rectangle Construction
We will now construct a specific family of dynamical rectangles that contain each neighborhood of $O_{r_{0}}$. Recall that by Lemma 12, for $Q_{0}=\left(\frac{1}{4},\frac{1}{4}\right)$, we have that
@@
for each $n,m\geq 0$ and $\lambda\in[0,\lambda_{2}]$, and as pointed out before Lemma 10, $\text{Orb}_{0}(Q_{0})$ is a finite set and bounded away from each $Q_{i}$ by at least $\frac{1}{4}$.
@@
This dynamical behavior will allow us to construct a Cantor set contained in $\mathcal{C}\left(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{O}_{r(\rho)}}\right)$ that has Hausdorff dimension greater than $\frac{\log 2}{\log\left(2+\frac{1}{M}\right)}$.
@@ 4.6.1 Base Set of Rectangles
Recalling the constant $C^{*}$ from (4.2) and $C^{**}$ from Lemma 10, we have (see Figure 9)
@@
    $\partial^{s}R_{m}(\lambda,Q_{i})\subset W^{s}_{\lambda}(Q_{0},m)$ and in particular, +$\partial^{s}_{+}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{+}(Q_{i})$ and $\partial^{s}_{-}R_{m}(\lambda,Q_{i})$ connects opposite boundaries of $\tilde{K}_{-}(Q_{i})$.
@@
    For any $P\in\partial^{s}R_{m}(\lambda,Q_{i})$, we have $|P-Q_{i}|<\frac{C^{*}}{2}$.
@@
We know that $W^{s}_{0}(Q_{0},m)$ is a line through $Q_{0}$ of slope $-\frac{1}{[a_{m},a_{m+1},\dots]}$ (see Lemma 3) and we know from Lemma 13 that
@@
Notice that $f_{0,\tilde{n},m}(P)\in W^{s}_{0,\text{loc}}(Q_{0},m+\tilde{n}+1)$, and so if $n\geq\tilde{n}$, then $f_{\lambda,n,m}(P)\in B_{\frac{1}{10}}(Q_{0})$ so that its distance from $\bigcup_{i=1}^{4}\{Q_{i}\}$ is greater than $\frac{3}{20}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=}\frac{1}{4}-\frac{1}{10}$.
@@
From Corollary 1 and Lemma 12, the quantity
$\sup\limits_{\begin{subarray}{c}k\in\mathbb{N}\\ 0\leq n\leq\tilde{n}\\ m\in\mathbb{N}\end{subarray}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{dist}_{H}}\left(f^{-1}_{\lambda,n,m}W^{s}_{\lambda,\text{loc}}(Q,k),f^{-1}_{0,n,m}W^{s}_{0,\text{loc}}(Q,k)\right)$ (4.16)